Goldbach's Conjecture Python

Loop through integers 4 through 100

for each number show two prime that sum up to integer

b)Show with  comments  how you found the prime numbers that add up to make each integer from the range 4 to 100

Example output:

4 = 2 + 2

6 = 3 + 3

etc.

Hint:You are finding the even numbers from 4 through a 100 and finding the two prime numbers that  add to make the integer and printing it out

like this

4 = 2 +2

6 = 3 + 3

8 = 3 + 5

and so on till you can gone through the range

Comment out code judiciously
Answer #1

#source code:

def pr(n): #it was used for the given number is prime or not

count=0

for i in range(1,n):

if(n%i==0):

count=count+1

if(count==1):

return True

return False

for k in range(4,100): #iterate 4 to 100

for i in range(2,100): #for first number

for j in range(2,100): #for second number

if(pr(i) and pr(j)):

if(i+j==k): #here made sum

print("{}={}+{}".format(k,i,j)) #here print the output

break

break

1 2 3 4 5 6 7 8 9 10 11 12 13 14 15 16 def pr(n): #it was used for the given number is prime or not count=0 for i in range(1,

#output:

user@user-Latitude-3490:-/Desktop$ python3.7 1.py 4=2+2 5=3+2 7=5+2 9=7+2 13=11+2 15=13+2 19=17+2 21=19+2 25=23+2 31=29+2 33=

#if you have any doubt or more information needed comment below..i will respond as possible as soon..thanks..